A candy bag contains red, orange, green, and blue sour balls. There are 12 red, 6 orange, 21 green, and 18 blue sour balls in the bag. What are the odds of reaching in the bag and randomly pulling out a sour ball that is neither red or green
Answer by htmentor(1343)

The total number of sour balls in the bag is 12+6+21+18 = 57.
The balls that are neither red or green are the blue and orange balls, of which there are 18 and 6, respectively.
Therefore P(neither red or green) = (18+6)/57 = 24/57 = 0.421